OpenStudy (arbyj):

helper will get a medal! 3x + 2y = A 5x + y = B Eliminating the variable y from the system of equations results in -7x = A - 2B 2B - A A + 2B

OpenStudy (photon336):

how do you think you would get rid of y?

OpenStudy (arbyj):

its been a while since ive done these problems uh... doesn't it have to be opposite of the other y to cancel it out?

OpenStudy (robtobey2):

\[\text{Eliminate}[\{3 x+2 y=A,5 x+y=B\},\{y\}] \]\[A+7 x=2 B \]-7x=A-2B

OpenStudy (legomyego180):

you still need help?

OpenStudy (arbyj):

@legomyego180 oh no, i got it, thank you! :)